OpenStudy (anonymous):

what is the given expression in 16^2 by 4^3?

OpenStudy (anonymous):

16^2*4^3

OpenStudy (anonymous):

by is like a square or rectangle so multiply it is that what you were looking for? :D hope I helped

ganeshie8 (ganeshie8):

may be its liek this ? \(\large \frac{16^2}{4^3}\)

OpenStudy (anonymous):

ok the original question is translate the phrase into a mathematical expression and then evaluate it. sixteen squared divided by four cubed. I got what I posted above. but im just confused about hat the expression is

OpenStudy (anonymous):

is what I posted the expression?

ganeshie8 (ganeshie8):

yea it makes sense now :) look belwo :- 16 squared = \(16 \times 16 = 16^2 \)

ganeshie8 (ganeshie8):

4 cubed = \(4 \times 4 \times 4 = 4^3\)

OpenStudy (anonymous):

16^2/4^3

OpenStudy (anonymous):

thats the problem in the words

OpenStudy (anonymous):

I have a lot to learn about this lol. thank you very much for your help.

OpenStudy (anonymous):

no problem :D